OpenStudy (photon336):

In any given production run, a manufacturer produces parts with defects at a rate of one out of every ten parts produced. If a customer randomly tests three parts off the production line, what is the probability that no more than one of the parts will have a defect?

OpenStudy (photon336):

Yeah, basically I know the probability of getting a defect is \[P(D)_{Probability~of~getting~a~defect} = \frac{ 1 }{ 10} \] so the complement of that, the probability a defect won't happen is \[1-P(D) = \frac{ 9 }{ 10 }\] If three parts are randomly selected. Trying to figure out what's the probability that no more than 1 of the parts will have a defect.

OpenStudy (photon336):

Was thinking that since we're looking for say 1/3 parts will have a defect. \[\frac{ number~of~desired~outcomes }{ total~possible~outcomes }\] ~\[\frac{ 1 }{ 3 }\] then I thought this was the chance we could get a part with a defect out of the three, but I knew that the chance had to be much lower than 1/10 so what I did was this. \[P(B)*P(D) = (\frac{ 1 }{ 10})*(\frac{ 1 }{ 3 }) = (\frac{ 1 }{ 30 })\]

OpenStudy (photon336):

@Michele_Laino are you good with probability?

OpenStudy (photon336):

@Owlcoffee

OpenStudy (photon336):

@mayankdevnani

OpenStudy (michele_laino):

no, I'm sorry, I'm not good with probability

OpenStudy (photon336):

@Kainui help? @TheSmartOne

OpenStudy (photon336):

@ayonnaleflore

OpenStudy (ayonnaleflore):

what exactly are you trying to find

OpenStudy (photon336):

If a customer randomly tests three parts off the production line, what is the probability that no more than one of the parts will have a defect?

OpenStudy (photon336):

basically you take 3 parts off the assembly line and say, well what's the probability one of those parts will be defective.

OpenStudy (ayonnaleflore):

it would be 1 out of 3

OpenStudy (ayonnaleflore):

simply because only one might be defective and there are three production lines

OpenStudy (ayonnaleflore):

@Photon336 what are the answer choices?

OpenStudy (photon336):

there aren't any.

OpenStudy (ayonnaleflore):

okay the answer would be \[\frac{ 3 }{ 9 }\] simplified to \[\frac{ 1 }{ 3 }\]

OpenStudy (mayankdevnani):

i'm not sure about the answer

OpenStudy (mayankdevnani):

@imqwerty @ParthKohli

Parth (parthkohli):

If none of them is to have a defect, then the probability is\[\left(\frac{9}{10}\right)^3\]If exactly one of them is to have a defect, then the probability is\[\binom{3}1\cdot \frac{1}{10} \cdot \left(\frac{9}{10}\right)^2\]...I think.

Parth (parthkohli):

So we just add those two, yeah

OpenStudy (mayankdevnani):

yep, i think too like this !

OpenStudy (mayankdevnani):

@Photon336 do you know the answer?

OpenStudy (agent0smith):

Just use binomial probability.

OpenStudy (agent0smith):

http://www.stat.yale.edu/Courses/1997-98/101/binprob.gif n=3, p=0.1 You want P(X=<1) and can use \[\large P(X \le 1) = P(X=0)+P(X=1)\]

OpenStudy (mayankdevnani):

Wow! i forgot this (my bad)

OpenStudy (mayankdevnani):

but what about PARTHKOHLI'S answer ?

OpenStudy (agent0smith):

Yes, that's correct.

OpenStudy (photon336):

The answer key says 97.2%

OpenStudy (agent0smith):

If you use binomial probability, you'll get that.

OpenStudy (photon336):

I didn't learn that yet I'll definitely look into that formula

OpenStudy (photon336):

My thinking was that, well if you had a 1/10th chance of getting a defective part, I was thinking to multiply that by (1/10)*(1/3) but it's wrong.

OpenStudy (agent0smith):

Are you sure you haven't learned it? Because otherwise you'd just be deriving the binomial probability formula yourself, to solve this.

OpenStudy (photon336):

so essentially there's no way to do this problem without binomial probability formula?

OpenStudy (agent0smith):

Yes. Because this is a binomial probability situation. What ParthKohli did was really the binomial formula.

OpenStudy (photon336):

I see thank you I'll have to look at this again